OpenStudy (anonymous):

somebody please help. i understand what it'd asking i just dont know how to set it up and solve it. Brian is creating a collage on a piece of cardboard that has an area of 110r3 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r to the fifth power square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (anonymous):

@splatter6

OpenStudy (anonymous):

i will help but first i have to finish this test so hang on

OpenStudy (anonymous):

okkkk

OpenStudy (anonymous):

did you try google because there is another few disscusion posts about this on openstudy

OpenStudy (anonymous):

yes i did but idk how they got the answers. and i copied it and pasted it the first time to my test and got caught so i want to actually know how to do it and explain it in a way i can understand

OpenStudy (anonymous):

Found this one online i'm sorry I can't provide you with an explanation or an answer since i don't know much about square roots outside of factoring them Since each piece of paper has the same area ( √r^5 cm^2 ) You need to put √r^5 into exponential notation --> √r^5 = r^5/2 This can be thought of in two ways: a) the square root of r taken to the fifth power or b) finding the square root of the fifth power or r. they are equivalent Multiply it by a variable and set it equal to 100r^3 cm^2. N(r^(5/2) ) =100 r^3 N = 100 r63 / (r^(5/2)) = 100 r^(3 - (5/2)) = 100r^(1/2) N = 100 r^(1/2) = 100√r

OpenStudy (anonymous):

yea i saw this one. thanks though

OpenStudy (anonymous):

ill still give u a medal

OpenStudy (anonymous):

i tried lol ur welcome try and get an admin on here just dont spam them with tags lol

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

its alright i dont mind but others get kinda mad

OpenStudy (anonymous):

yea like yest

OpenStudy (anonymous):

yup lol

OpenStudy (anonymous):

(: lol

OpenStudy (anonymous):

try bumping your question to the top

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

yup and now i have to wait 3:45

OpenStudy (anonymous):

would u be able to help me with 2-3 more different questions

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is the simplified form of